A ball rolls up in an inclined plane with an initial angular velocity 100



(a) if its radius is r. What is it initial linear velocity? After the time t the ball stops.



(b) what is it angular retardation



(C) what is the angular displacement before its stops

1
Expert's answer
2022-02-06T14:34:11-0500

Explanations and Calculations


a)

  • Initial linear velocity is "\small V=r\omega=100r"


b)

  • If the angular retardation is "\small \alpha" ,

"\qquad\qquad
\begin{aligned}
\small \omega_f&=\small \omega_i+\alpha.t\\
\small \alpha&=\small \frac{0-100}{t}\\
&=\small -\frac{100}{t}
\end{aligned}"

c)

  • If the angular displacement is "\small \theta,"

"\qquad\qquad
\begin{aligned}
\small \theta&=\small \omega.t+\frac{1}{2} \alpha t^2\\
&=\small 100t-\frac{1}{2}\times\frac{100}{t}\times t^2\\
&=\small 100t-50t\\
&=\small 50t
\end{aligned}"